# A low-rank solver for conforming multipatch Isogeometric Analysis

### v3

@@ 3.4 Subdomains as unions of patches
We consider the following low-rank approximations of the entries of $Q^{(i,k,\ell)}$:
$Q^{(i,k,\ell)}_{\alpha,\beta}\left(\xi_{1},\xi_{2},\xi_{3}\right)\approx\sum_{r_{1},r_{2},r_{3}}\mathfrak{q}_{r_{1},r_{2},r_{3}}^{(i,k,\ell,\alpha,\beta)}\cdot q^{(i,k,\ell,\alpha,\beta)}_{(1,r_{1})}\left(\xi_{1}\right)\cdot q^{(i,k,\ell,\alpha,\beta)}_{(2,r_{2})}\left(\xi_{2}\right)\cdot q^{(i,k,\ell,\alpha,\beta)}_{(3,r_{3})}\left(\xi_{3}\right).$ (3.16)
This approximation can be computed using the same strategy as for (3.14). Plugging (3.16) into (3.15), and exploiting the tensor structure of $\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}$, we obtain the Tucker approximation (3.5).
@@
Since the low-rank strategy is meant to effectively tackle large problems, we have considered above the subdomains as union of two adjacent patches. Other strategies are possible, resulting e.g. in a larger number of smaller subdomains. We emphasize that a large number of subdomains would require the introduction of a coarse solver. Indeed, it is well-known that preconditioners based on local domain decomposition without coarse solver are not robust with respect to the number of subdomains.
@@ 4 Numerical experiments
In all the tests, we employ a dyadic refinement and every patch has the same number of elements $n_{el}=2^{L}$ in each parametric direction and for different discretization levels $L$. The numerical solution is computed with the TPCG method, discussed in Section 3.3, with relative tolerance $tol=10^{-6}$ and the zero vector as initial guess. The system matrix and the right-hand side are approximated with the technique of Section 3.4 by setting the relative tolerance of Chebfun3F equal to $10^{-1}tol=10^{-7}$ (see [8] for more details). The dynamic truncation described in Section 3.3 is used fixing the initial relative tolerance as $\epsilon_{0}=10^{-1}$ and the minimum relative tolerance as $\epsilon_{\min}=tol\|\boldsymbol{f}\|_{2}10^{-1}$. The threshold $\delta$ is set equal to $10^{-3}$, the reducing factor $\alpha$ is set equal to 0.5. The relative tolerance parameter $\beta$ is chosen as $10^{-1}$, while the intermediate relative tolerance $\gamma$ is chosen as $10^{-8}$. The relative tolerance $\epsilon_{rel}^{prec}$ for the approximation of the diagonal blocks of the preconditioner (see [21] for details) is set equal to $10^{-1}$. We experimentally verified that these choices of the parameters yield good performance of the TPCG method.
@@ 4.1 L-shaped domain
In this test we consider a three dimensional L-shaped domain, as represented in Figure LABEL:fig:in_dom, where different colors represent different patches. We have two subdomains, one is the union of the yellow and light green patches and the other the union of the light green and dark green patches.
For this domain, we preliminary perform a convergence test to assess that the numerical solution yielded by our approach is indeed a good approximation of the Galerkin solution. For this purpose, we consider a Poisson toy problem with homogeneous Dirichlet boundary conditions and source function $f(x,y,z)=\sin(2\pi x)\sin(2\pi y)\sin(2\pi z)$. Note that the exact solution of this toy problem is $u(x,y,z)=-\left(12\pi^{2}\right)^{-1}f(x,y,z)$. We assemble and solve the problem with a low-rank strategy analogous to that described previously (in particular we fix $10^{-10}$ as tolerance for TPCG) and then compute the $H^{1}$ and $L^{2}$ norms of the error for the obtained solution, for different discretization levels and polynomial degrees. Results are shown in Figure 1 and confirm the validity of our approach.
Figure 1: Errors for a Poisson problem in the L-shaped domain.
[…]
Figure 2: Initial geometry and numerical solution for the L-shaped domain.
@@
In Table 1 we report the number of iterations for degrees $p=2,3,4,5$ and number of elements per patch per parametric direction $n_{el}=2^{L}$ with $L=5,6,7,8.$ The number of iterations is almost independent of the degree and grows only mildly with the number of elements $n_{el}$. We emphasize that for the finer discretization level the number of degrees of freedom is roughly 200 millions.
@@
[…]
@@
Figure 4: Initial geometry and numerical solution for the 3D-cross domain.
The number of iterations for degrees $p=2,3,4,5$ and number of elements per parametric direction $n_{el}=2^{L}$ with $L=5,6,7,8$ is reported in Table 2. The number of iterations is almost independent on the degree and mesh-size. These numbers are similar, although slightly higher, to the ones of Table 1 relative to the L-shaped domain: the proposed low-rank strategy seems to be independent on the number of subdomains to which a patch belongs.
@@
4.3 Ring domain
[…]
Figure 6: Initial geometry and numerical solution for the ring domain.

### v4

@@ 3.4 Subdomains as unions of patches
The low-rank approximations of the entries of $Q^{(i,k,\ell)}$ are computed similarly as for (3.18). We assume as before that $\Theta^{(i)}=\Omega^{(i_{1})}\cup\Omega^{(i_{2})}$. For each $\alpha,\beta=1,2,3,$ we use again the approximation (3.16) of the coefficients $C_{\alpha,\beta}^{i_{1},k,\ell}$ and $C_{\alpha,\beta}^{i_{2},k,\ell}$. The two approximations are combined as in (3.20) and the corresponding low-rank representation is truncated with a tolerance close to machine precision to remove any redundancy. In the end, the approximations of the entries of $Q^{(i,k,\ell)}$ are then plugged into (3.22), and thanks to the tensor structure of $\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}$, we obtain the Tucker approximation (3.6). As it will be useful in the following, we observe that this approach yields
$\boldsymbol{v}^{T}\widetilde{\mathbf{A}}^{(i,i,k,\ell)}\boldsymbol{w}=\int_{[0,1]^{3}}\left(\nabla\widehat{v}_{ptc}^{(i_{i})}\right)^{T}\widetilde{C}^{(i_{1},k,\ell)}\;\nabla\widehat{w}_{ptc}^{(i_{1})}\;+\;\left(\nabla\widehat{v}_{ptc}^{(i_{2})}\right)^{T}\widetilde{C}^{(i_{2},k,\ell)}\;\nabla\widehat{w}_{ptc}^{(i_{2})}\;\mathrm{d}{\underline{\xi}},$ (3.23)
where $\widehat{v}_{ptc}^{(m)}=v_{h}\circ\mathcal{F}_{m}$ and $\widehat{w}_{ptc}^{(m)}=w_{h}\circ\mathcal{F}_{m}$, for $m=i_{1},i_{2}$.
@@
Since the low-rank strategy is meant to effectively tackle large problems, we have considered above the subdomains as union of adjacent patches. It would be of course possible to consider a larger number of smaller subdomains. We emphasize, however, that a large number of subdomains would require the introduction of a coarse solver. Indeed, it is well-known that preconditioners based on local domain decomposition without coarse solver are not robust with respect to the number of subdomains.
4 Analysis of the low-rank approximation of the linear system
We now show some theoretical results related to the low-rank approximation of the linear system (3.8). We adopt the following notation: given $\underline{v}_{h}\in\left[V_{h}\right]^{3}$, we denote with $\nabla\underline{v}_{h}$ the vector of $\mathbb{R}^{9}$ obtained by stacking the gradients of the components of $\underline{v}_{h}$.
Let $\underline{v}_{h},\underline{w}_{h}\in\left[V_{h}\right]^{3}$, and let $\boldsymbol{v},\boldsymbol{w}$ be block vectors representing these functions with respect to $\left[V_{h}^{(1)}\right]^{3},\ldots,$
$\left[V_{h}^{(\mathcal{N}_{sub})}\right]^{3}$ and their respective bases. We preliminary observe that
$\boldsymbol{v}^{T}\mathbf{A}\boldsymbol{w}=\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\nabla\widehat{\underline{v}}_{ptc}^{(m)}\right)^{T}C^{(m)}\;\nabla\widehat{\underline{w}}_{ptc}^{(m)}\;\mathrm{d}{\underline{\xi}},$ (4.1)
where $\widehat{\underline{v}}_{ptc}^{(m)}=\underline{v}_{h}\circ\mathcal{F}_{m}$, $\widehat{\underline{w}}_{ptc}^{(m)}=\underline{w}_{h}\circ\mathcal{F}_{m}$, and
$C^{(m)}(\underline{\xi})=\begin{bmatrix}C^{(m,1,1)}(\underline{\xi})&C^{(m,1,2)}(\underline{\xi})&C^{(m,1,3)}(\underline{\xi})\\ C^{(m,2,1)}(\underline{\xi})&C^{(m,2,2)}(\underline{\xi})&C^{(m,2,3)}(\underline{\xi})\\ C^{(m,3,1)}(\underline{\xi})&C^{(m,3,2)}(\underline{\xi})&C^{(m,3,3)}(\underline{\xi})\end{bmatrix}\in\mathbb{R}^{9\times 9},$ (4.2)
whose blocks are defined in (3.13).
In the next lemma we show in particular that $\widetilde{\mathbf{A}}$ defines a bilinear form analagous to (4.1).
Lemma 1.
Let $\underline{v}_{h},\underline{w}_{h}\in\left[V_{h}\right]^{3}$, and let $\boldsymbol{v},\boldsymbol{w}$ be block vectors representing these functions with respect to $\left[V_{h}^{(1)}\right]^{3},\ldots,\left[V_{h}^{(\mathcal{N}_{sub})}\right]^{3}$ and their respective bases. Then
$\boldsymbol{v}^{T}\widetilde{\mathbf{A}}\boldsymbol{w}$ $=\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\nabla\widehat{\underline{v}}_{ptc}^{(m)}\right)^{T}\widetilde{C}^{(m)}\;\nabla\widehat{\underline{w}}_{ptc}^{(m)}\;\mathrm{d}{\underline{\xi}},$ (4.3)
$\boldsymbol{v}^{T}\widetilde{\boldsymbol{f}}$ $=\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\widehat{\underline{v}}_{ptc}^{(m)}\right)^{T}\underline{\widetilde{g}}^{(m)}\;\mathrm{d}{\underline{\xi}},$ (4.4)
where $\widehat{\underline{v}}_{ptc}^{(m)}=\underline{v}_{h}\circ\mathcal{F}_{m}$, $\widehat{\underline{w}}_{ptc}^{(m)}=\underline{w}_{h}\circ\mathcal{F}_{m}$, $\widetilde{C}^{(m)}(\underline{\xi})$ is the $3\times 3$ block matrix whose blocks are the $3\times 3$ matrices $\widetilde{C}^{(m,k,\ell)}(\underline{\xi})$ from (3.14) for $k,\ell=1,2,3$, while $\underline{\widetilde{g}}^{(m)}(\underline{\xi})=\begin{bmatrix}\widetilde{g}_{1}^{(m)}(\underline{\xi})&\widetilde{g}_{2}^{(m)}(\underline{\xi})&\widetilde{g}_{3}^{(m)}(\underline{\xi})\end{bmatrix}^{T}$, whose entries are introduced in (3.19).
Proof.
By considering (3.15) and (3.23) for $k,\ell=1,2,3,$ we infer
$\boldsymbol{v}^{T}\widetilde{\mathbf{A}}\boldsymbol{w}=\sum_{i,j=1}^{\mathcal{N}_{sub}}\left(\boldsymbol{v}^{(i)}\right)^{T}\widetilde{\mathbf{A}}^{(i,j)}\boldsymbol{w}^{(j)}=\sum_{i,j=1}^{\mathcal{N}_{sub}}\;\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\nabla\widehat{\underline{v}}_{ptc}^{(m,i)}\right)^{T}\widetilde{C}^{(m)}\nabla\widehat{\underline{w}}_{ptc}^{(m,j)}\;\mathrm{d}{\underline{\xi}},$
where $\widehat{\underline{v}}_{ptc}^{(m,i)}=\underline{v}_{h}^{(i)}\circ\mathcal{F}_{m}$ and $\widehat{\underline{v}}_{ptc}^{(m,j)}=\underline{v}_{h}^{(j)}\circ\mathcal{F}_{m}$. Note that at least one between $\widehat{\underline{v}}_{ptc}^{(m,i)}$ and $\widehat{\underline{w}}_{ptc}^{(m,j)}$ is the null function unless $\Omega^{(m)}\subseteq\Theta^{(i)}\cap\Theta^{(j)}$. By switching the order of the sums, we obtain
$\sum_{m=1}^{\mathcal{N}_{ptc}}\sum_{i,j=1}^{\mathcal{N}_{sub}}\int_{[0,1]^{3}}\left(\nabla\widehat{\underline{v}}_{ptc}^{(m,i)}\right)^{T}\widetilde{C}^{(m)}\nabla\widehat{\underline{w}}_{ptc}^{(m,j)}\;\mathrm{d}{\underline{\xi}}=\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\sum_{i=1}^{\mathcal{N}_{sub}}\nabla\widehat{\underline{v}}_{ptc}^{(m,i)}\right)^{T}\widetilde{C}^{(m)}\left(\sum_{j=1}^{\mathcal{N}_{sub}}\nabla\widehat{\underline{w}}_{ptc}^{(m,j)}\right)\;\mathrm{d}{\underline{\xi}}.$
Then (4.3) follows by observing that $\sum_{i=1}^{\mathcal{N}_{sub}}\widehat{\underline{v}}_{ptc}^{(m,i)}=\widehat{\underline{v}}_{ptc}^{(m)}$ and $\sum_{j=1}^{\mathcal{N}_{sub}}\widehat{\underline{w}}_{ptc}^{(m,j)}=\widehat{\underline{w}}_{ptc}^{(m)}$. Similarly, considering (3.21) for $k=1,2,3$, we infer
$\boldsymbol{v}^{T}\widetilde{\boldsymbol{f}}=\sum_{i=1}^{\mathcal{N}_{sub}}\left(\boldsymbol{v}^{(i)}\right)^{T}\widetilde{\boldsymbol{f}}^{(i)}=\sum_{i=1}^{\mathcal{N}_{sub}}\sum_{i=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\widehat{\underline{v}}_{ptc}^{(m,i)}\right)^{T}\underline{\widetilde{g}}^{(m)}\;\mathrm{d}{\underline{\xi}}=\sum_{i=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\sum_{i=1}^{\mathcal{N}_{sub}}\widehat{\underline{v}}_{ptc}^{(m,i)}\right)^{T}\underline{\widetilde{g}}^{(m)}\;\mathrm{d}{\underline{\xi}}$
which proves (4.4). ∎
The following result shows that the low-rank approximation described in Section 3.4 does not alter the kernel of the system matrix, and that problem (3.8) admits a solution. This result is based on the assumption that the matrix $\widetilde{C}^{(m)}(\underline{\xi})$ introduced in Lemma 1 is positive definite everywhere, which is clearly satisfied provided that the low-rank approximations (3.14) are accurate enough.
Theorem 1.
Assume that the matrix $\widetilde{C}^{(m)}(\underline{\xi})$ defined in Lemma 1 is positive definite for every $m=1,\ldots,\mathcal{N}_{ptc}$ and for every $\underline{\xi}\in[0,1]^{3}$. Then it holds
$\ker\left(\widetilde{\mathbf{A}}\right)=\ker\left(\mathbf{A}\right)\qquad\text{and}\qquad\widetilde{\boldsymbol{f}}\in\mathrm{range}\left(\widetilde{\mathbf{A}}\right).$
Proof.
Let $\boldsymbol{v}$ be a block vector representing the function $\underline{v}_{h}\in\left[V_{h}\right]^{3}$. Thanks to Lemma 1, we have
$\boldsymbol{v}^{T}\widetilde{\mathbf{A}}\boldsymbol{v}=\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\nabla\widehat{\underline{v}}_{ptc}^{(m)}\right)^{T}\widetilde{C}^{(m)}\nabla\widehat{\underline{v}}_{ptc}^{(m)}\;\mathrm{d}{\underline{\xi}}\geq\min_{m=1,\ldots,\mathcal{N}_{ptc}}\;\min_{\underline{\xi}\in[0,1]^{3}}\lambda_{\min}\left(\widetilde{C}^{(m)}(\underline{\xi})\right)\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\|\nabla\widehat{\underline{v}}_{ptc}^{(m)}\|_{2}^{2}\;\mathrm{d}{\underline{\xi}}.$
The latter term is zero if and only if $\widehat{\underline{v}}_{ptc}^{(m)}$ is the null function for every $m=1,\ldots,\mathcal{N}_{ptc}$, which is equivalent to require that $\underline{v}_{h}$ is the null function. Since $\ker\left(\mathbf{A}\right)$ is the set of vectors representing the null function, we have that $\ker\left(\mathbf{A}\right)=\ker\left(\widetilde{\mathbf{A}}\right)$. Then, from (4.4) we infer that $\boldsymbol{v}^{T}\widetilde{\boldsymbol{f}}=0$ whenever $\boldsymbol{v}$ is a block vector representing the null function. Therefore, $\widetilde{\boldsymbol{f}}\in\ker\left(\widetilde{\mathbf{A}}\right)^{\perp}=\text{range}\left(\widetilde{\mathbf{A}}\right)$.
We are now ready to prove an upper bound on the relative error in the 2-norm committed by approximating $\mathbf{A}$ with $\widetilde{\mathbf{A}}$, in terms of the accuracy of the low-rank approximations (3.14).
Theorem 2.
Assume that, for every $\alpha,\beta,k,\ell=1,2,3$ and for every $m=1,\ldots,\mathcal{N}_{ptc}$, the low-rank approximations (3.14) satify
$\max_{\underline{\xi}\in[0,1]^{3}}|C_{\alpha,\beta}^{(m,k,\ell)}(\underline{\xi})-\widetilde{C}_{\alpha,\beta}^{(m,k,\ell)}(\underline{\xi})|\leq\zeta$ (4.5)
for some $\zeta\geq 0$. Then
$\frac{\|\mathbf{A}-\widetilde{\mathbf{A}}\|_{2}}{\|\mathbf{A}\|_{2}}\leq\frac{9\;\zeta}{\displaystyle\min_{m=1,\ldots,\mathcal{N}_{ptc}}\;\min_{\underline{\xi}\in[0,1]^{3}}\lambda_{\min}\left(C^{(m)}\left(\underline{\xi}\right)\right)},$
where $\lambda_{\min}\left(C^{(m)}\left(\underline{\xi}\right)\right)$ is the minimum eigenvalue of the matrix $C^{(m)}\left(\underline{\xi}\right)\in\mathbb{R}^{9\times 9}$ defined in (4.2).
Proof.
Since $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ are symmetric, it holds
$\|\mathbf{A}-\widetilde{\mathbf{A}}\|_{2}=\max_{\boldsymbol{v}}\frac{\boldsymbol{v}^{T}\left(\mathbf{A}-\widetilde{\mathbf{A}}\right)\boldsymbol{v}}{\|\boldsymbol{v}\|_{2}^{2}}.$
Exploiting (3.15) and (3.23), we infer
$\boldsymbol{v}^{T}\left(\mathbf{A}-\widetilde{\mathbf{A}}\right)\boldsymbol{v}$ $=\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\nabla\widehat{\underline{v}}_{ptc}^{(m)}\right)^{T}\left(C^{(m)}-\widetilde{C}^{(m)}\right)\nabla\widehat{\underline{v}}_{ptc}^{(m)}\;\mathrm{d}{\underline{\xi}}$
$\leq\max_{m=1,\ldots,\mathcal{N}_{ptc}}\;\max_{\underline{\xi}\in[0,1]^{3}}\lambda_{\max}\left(C^{(m)}(\underline{\xi})-\widetilde{C}^{(m)}(\underline{\xi})\right)\;\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\|\nabla\widehat{\underline{v}}_{ptc}^{(m)}\|^{2}_{2}\;\mathrm{d}{\underline{\xi}}$
$\leq\frac{\displaystyle\max_{m=1,\ldots,\mathcal{N}_{ptc}}\;\max_{\underline{\xi}\in[0,1]^{3}}\lambda_{\max}\left(C^{(m)}(\underline{\xi})-\widetilde{C}^{(m)}(\underline{\xi})\right)}{\displaystyle\min_{m=1,\ldots,\mathcal{N}_{ptc}}\;\min_{\underline{\xi}\in[0,1]^{3}}\lambda_{\min}\left(C^{(m)}\left(\underline{\xi}\right)\right)}\sum_{m=1}^{\mathcal{N}_{ptc}}\int_{[0,1]^{3}}\left(\nabla\widehat{\underline{v}}_{ptc}^{(m)}\right)^{T}C^{(m)}\nabla\widehat{\underline{v}}_{ptc}^{(m)}\;\mathrm{d}{\underline{\xi}}$
$=\frac{\displaystyle\max_{m=1,\ldots,\mathcal{N}_{ptc}}\;\max_{\underline{\xi}\in[0,1]^{3}}\lambda_{\max}\left(C^{(m)}(\underline{\xi})-\widetilde{C}^{(m)}(\underline{\xi})\right)}{\displaystyle\min_{m=1,\ldots,\mathcal{N}_{ptc}}\;\min_{\underline{\xi}\in[0,1]^{3}}\lambda_{\min}\left(C^{(m)}\left(\underline{\xi}\right)\right)}\;\boldsymbol{v}^{T}\mathbf{A}\boldsymbol{v}.$
Inequality (4.5) then follows by observing that $\max_{\boldsymbol{v}}\frac{\boldsymbol{v}^{T}\mathbf{A}\boldsymbol{v}}{\|\boldsymbol{v}\|_{2}^{2}}=\|\mathbf{A}\|_{2}$ and that
$\lambda_{\max}\left(C^{(m)}(\underline{\xi})-\widetilde{C}^{(m)}(\underline{\xi})\right)\leq\|C^{(m)}(\underline{\xi})-\widetilde{C}^{(m)}(\underline{\xi})\|_{\infty}\leq 9\;\max_{\alpha,\beta,k,\ell=1,2,3}|C_{\alpha,\beta}^{(m,k,\ell)}(\underline{\xi})-\widetilde{C}_{\alpha,\beta}^{(m,k,\ell)}(\underline{\xi})|\leq 9\;\zeta$
for every $m=1,\ldots,\mathcal{N}_{ptc}$ and for every $\underline{\xi}\in[0,1]^{3}$.
@@ 5 Numerical experiments
[…]
@@ 5.1 L-shaped domain
In this test we consider a three dimensional L-shaped domain, as represented in Figure 2, where different colors represent different patches. We have two subdomains, one is the union of the yellow and light green patches and the other the union of the light green and dark green patches. We consider homogeneous Dirichlet boundary conditions on the patch faces that lie on the planes $\{x=-1\}\cup\{z=1\}\cup\{x=0\}\cup\{z=0\}$, and homogeneous Neumann boundary conditions elsewhere.
Figure 2: L-shaped domain.
@@
In Table 1 we report the number of iterations for degrees $p=3,4,5$ and number of elements per patch per parametric direction $n_{el}=2^{L}$ with $L=5,6,7,8.$ The number of iterations is almost independent of the degree and grows only mildly with the number of elements $n_{el}$. We emphasize that for the finer discretization level the number of degrees of freedom is roughly 200 millions.
@@
In this test, we consider a 3D-cross domain, as represented in Figure 4. Note that in this domain we have that the central cube belongs to six subdomains, which is the maximum allowed by our strategy. We consider homogeneous Dirichlet boundary conditions everywhere, except for the external faces. The multilinear ranks of the blocks of $\widetilde{\mathbf{A}}$ are
@@
Figure 4: 3D-cross domain.
The number of iterations for degrees $p=3,4,5$ and number of elements per parametric direction $n_{el}=2^{L}$ with $L=5,6,7,8$ is reported in Table 2. The number of iterations is almost independent on the degree and mesh-size. These numbers are similar, although slightly higher, to the ones of Table 1 relative to the L-shaped domain: the proposed low-rank strategy seems to be independent on the number of subdomains to which a patch belongs.
@@
5.3 Thick square with jumping coefficient
In this test, we consider a thick square $[0,2]\times[0,2]\times[0,1]$ composed by 9 patches, as represented in Figure 6. According to the procedure described at the beginning of Section 3.4, we split $\Omega$ into 4 subdomains, each obtained by merging the 4 patches around an inner edge. We impose homogeneous Dirichlet boundary conditions everywhere, except for the side $\{z=1\}$, where we impose homogeneous Neumann boundary conditions. We consider jumping Lamé parameters, by setting $\nu=0.3$ and $E=1$ on green patches and $E=6$ on red patches. The multilinear ranks of the blocks of $\widetilde{\mathbf{A}}$ are
$\left(R^{A}_{1}(i,i,k,k),R^{A}_{2}(i,i,k,k),R^{A}_{3}(i,i,k,k)\right)=(6,6,6),\qquad k=1,2,3,$
and
$\left(R^{A}_{1}(i,i,k,\ell),R^{A}_{2}(i,i,k,\ell),R^{A}_{3}(i,i,k,\ell)\right)=(4,4,4),\qquad k,\ell=1,2,3,\;k\neq\ell,$
for $i=1,\ldots,4$,
$\left(R^{A}_{1}(i,j,k,k),R^{A}_{2}(i,j,k,k),R^{A}_{3}(i,j,k,k)\right)=(3,3,3),\qquad k,\ell=1,2,3,\;k\neq\ell\;i\neq j,$
and
$\left(R^{A}_{1}(i,j,k,\ell),R^{A}_{2}(i,j,k,\ell),R^{A}_{3}(i,j,k,\ell)\right)=(2,2,2),\qquad k,\ell=1,2,3,\;k\neq\ell\;i\neq j,$
for $i,j=1,\ldots,4$.
Figure 6: Thick square domain.
In Table 3 we report the numbers of iterations for degrees $p=3,4,5$ and number of elements per direction $n_{el}=2^{L}$ with $L=4,5,6,7$. The number of iterations is almost independent of degree and mesh-size.
 Iteration number
$n_{el}$ $p=3$ $p=4$ $p=5$
16 68 65 66
32 71 67 69
64 71 70 72
128 78 73 75
Table 3: Number of iterations for the thick square domain.
The maximum of the multilinear ranks of the solution in all the directions and for all the subdomains is shownn in Figure LABEL:fig:rank_thick_square. As the number of elements per direction $n_{el}$ grows, the multilinear rank of the solution increases, but, as represented in Figure LABEL:fig:mem_thick_square, the memory storage for the solution hugely reduces with respect to the full solution.
Figure 7: Results for the thick square domain.
5.4 Thick ring domain
In this test we consider a thick ring domain, represented in Figure 9. We have four patches and four subdomains, represented by the union of green and yellow patches, yellow and blue patches, blue and red patches and red and green patches.
For this domain, we preliminary perform a convergence test to assess that the numerical solution yielded by our approach is indeed a good approximation of the Galerkin solution. For this purpose, we consider a Poisson toy problem with homogeneous Dirichlet boundary conditions and source function $f(x,y,z)=\sin(4\pi x)\sin(4\pi y)\sin(4\pi z)$. Note that the exact solution of this toy problem is $u(x,y,z)=-\left(48\pi^{2}\right)^{-1}f(x,y,z)$. We solve the problem with a low-rank strategy analogous to that described previously (in particular we fix $10^{-10}$ as tolerance for TPCG) and then compute the $H^{1}$ and $L^{2}$ norms of the error for the obtained solution, for different discretization levels and polynomial degrees. Results are shown in Figure 8 and confirm the validity of our approach.
Figure 8: Errors for a Poisson problem in the thick ring domain.
We now solve a linear elasticity problem on the same domain. We impose homogeneous Dirichlet boundary conditions on $\{z=0\}$ and in the central walls of the thick ring, while homogeneous Neumann boundary conditions are imposed elsewhere.
Figure 9: Thick ring domain.